1st year Course Details 2000-2001

Course  Code  M Tu W Th F  Start  First Meet  Half Comments
AI1Ah+Bh  AI0001+AI0002  4  4  -  4 -  Oct9+Jan29 LTh4/AT  1+2
Ast1Ah  PH0006  2  -  -  - 2+4 Oct9  LTh5/AT  1 Tut M3/F3
Ast1Bh  PH0007  2+4 -  -  - 2+4 Jan29  LTh1/AT  2 Tut M3/F3
Ast1Ch  PH0008  2+4 -  -  - 2  Oct9  LTh3/AT  1 Tut M3/F3
BusSt1  BS0001  5  5  -  5 -  Oct9  GeoSqLTh  1+2 +tut. CS/ManSc only.
CL1h  CS0001  5.10 5.10 - 5.10 -  Oct9  LTh4/AT  1
CS1Ah  CS0077  2  11 - 11 -  Oct9  LTh4/AT  1 +tut.
CS1Bh  CS0078  2  11 - 11 -  Jan29  LTh4/AT  1 +tut.
EE1h  EE0001  10  -  - 10 10 Jan29  LTh4/AT  2 + M/Tu & Th 2-5
Elect1h  EE0002  5  5  -  5 -  Jan29  LThB/JCMB 2 Lab pm as arranged
Eng1h  CE0021  10 10 10 10 10 Oct9  LTh4/AT  1 Tut. Tu/Th 2-5
HistSci1h  SU0003  5.10 5.10 - 5.10 -  Oct9  LTh1/AT  1
HumComm1h  CG0001  5.10 5.10 - 5.10 -  Jan29  LTh2/AT  2
IndM1h  BS0002  9  -  9  - 9  Oct9  LTh1/Sand.  1 +tut. Not with BusSt1. Sanderson Blg
Logic 1h  PP9976  5  5  -  5 -  Oct9  LTh3/AT  1
Ling1  LI0001  10 10 10 10 10 Oct9  LThB/DHT 1+2
Maths Hons 1  MS0087-90  12 12 12 12 12 Oct9  LTh2/AT  1+2 enrol for each quarter course
Maths1Y1h+2h MS0001+MS0003  12 12 12 12 12 Oct9(both) LTh3/AT  1+2 +tut - norm Fri12, thin
Other Maths 1  by module  very complicated
Meteor1  MY0001  10  - 10 - 10 Oct9  LThC/JCMB 1+2 Lab/Tut M2-3.30/Tu10-11.30/Th2-3.30
PhilSci1h  PP9977  5  5  -  5 -  Jan29  LTh3/AT  2
Physics1Ah+1Bh PH0088 + PH0089 11  - 11 - 11 Oct9+Jan29 LTh4/AT  1+2 Lab pm as arranged
Int2Phil1  PP0002  10  - 10 - 10 Oct9  LThA/DHT 1+2 Quota
Psych1  PS0001  11 11 11 11 11 Oct9  LThA/DHT 1+2 Quota
SciInSoc1Ah  SU0001  5  5  -  5 -  Jan29  LTh1/AT  2
StatSoc1h  MS0113  10 10 - 10 -  Oct9  LTh2/AT  1 tut norm W10
TMSE1h  ME0002  9  -  9  - 9  Jan29  LTh2/HB  2 Hudson Beare Blg

2nd year Course Details 2000-2001

Course Code M Tu W Th F Start Venue Half Comments
AI2Ah+Bh  AI0003+0004 - 9 - 9 9 Oct10+Jan30 LTh4AT (first half),D/FH (second half) 1+2

Business Studies 2 half-courses:

MS&IS  BS0035  5 5 - 5 - Oct9  G.08/WRB  1 +tut.
OMSC  BS0033  5 5 - 5 - Jan 29  F.21Psych  2 +tut. *** Take 2nd year
MS&OP  BS0040  5 5 - 5 - Jan29  G19/AFB  2 +tut. Normally 3rd year
CS2  CS0006+0007 11 - 11 - 11 Oct9+Jan29 LThA/JCMB (both)  1+2
EE2  EE0005  12 12 10 - 10 Oct9  LTh2/Hudson Beare  1+2 Tuts Th9+12. Lab M/F2-5
Ling2  LI0002  10 10 10 10 10 Oct10  G10 AFB  1+2
Maths Hons 2 by module  varies by module  Oct9  LThB/JCMB  1+2 complicated
Maths2Y  MS0014  12 12 12 12 12 Oct9  LThA/JCMB  1+2
Other Maths 2 by module  very complicated
Phil2Eh  PP9978  - 2 - 2 2 Jan30  3.01/DHT  2 “Logic 2h”
Psych 2  PS0002  11 11 - 11 - Oct9  G.08/WRB  1+2 Prac for 3hrs one afternoon (not Fri)